OpenStudy (anonymous):

Use substitution to solve. 3x+24y=24 27x-15y=-15

OpenStudy (anonymous):

multiply the top equation by -9 to cancel out the x's. Then you can solve for y. Once you have that, plug the y back in to find the x.

OpenStudy (zzr0ck3r):

this is not substitution @kjmoney

OpenStudy (zzr0ck3r):

solve for x in the first equation(you could solve for any variable in either equation but this one looks like the best choice) 3x+24y = 24 3x = 24-24y x = 8 - 8y now put this into equation2 for x 27x-15y = -15 so 27(8-8y) - 15y = -15 can you solve for y here @blkbutterfly13 ?

OpenStudy (anonymous):

@zzr0ck3r thanks, that's how I started off but it didn't look right. I just needed a second opinion.

OpenStudy (zzr0ck3r):

you got it now? solve for y, then plug that value into either equation and solve for x

OpenStudy (anonymous):

yes I did, thank you so much!!

OpenStudy (zzr0ck3r):

np